OpenStudy (anonymous):

what is the equation in point-slope form of the line passing through (3, 6) an (-2, 1)?

OpenStudy (matt101):

You want the equation to be written as y = mx + b, where (x,y) are the coordinates of any point on the line, m is the slope, and b is the y-intercept. You have x and y already from either point given in the question. To find the slope, find the change in rise (y) over the change in run (x): m = (6-1)/(3-(-2)) = 5/5 = 1 You have x, y, and m, so plug all three in to y = mx + b to find b: 6 = (1)(3) + b --> b = 3 Therefore your general equation is y = x + 3

OpenStudy (anonymous):

Thank you:) That's what I did in the beginning but y = x + 3...isn't that slope-intercept form? For my answer in point-slope form, I got y - 6 = -1(x + 3) but I think I did something wrong. Ahhh this sucks :( lol. I appreciate you helping me! I'd give you a medal on here, but I'm new to this site and I have no idea how to do that ):

OpenStudy (matt101):

Sorry yes for the point-slope formula, it's the same as the formula for slope. You've just multiplied both sides by x1-x2. The general formula is as follows: y-y1 = m(x-x1) Plugging in the numbers from above, we get y - 6 = x - 3 for the general form of the equation. You can see that this simplifies to y = x + 3, the slope-intercept form of the equation. As for medals, just click best answer :)

OpenStudy (anonymous):

Oh yayy :) Thank you so much!!